OpenStudy (rubylovesjaune):

Simplify. ^5 (square root) x^16

OpenStudy (legomyego180):

\[\sqrt[5]{x^16}\] ?

OpenStudy (rubylovesjaune):

Yes

OpenStudy (rubylovesjaune):

It has to be simplified from that thought, and I don't know how.

OpenStudy (anonymous):

so than ,,yes" this mean that not is squarroot - ok. ?

OpenStudy (anonymous):

looke please sqrtx= x^(1/2) can using this ?

OpenStudy (anonymous):

remember always the denominator of fractional exponent will meaning the index of radical ok. ?

OpenStudy (rubylovesjaune):

Okay

OpenStudy (anonymous):

so than use these in case of your exercise so what will get ?

OpenStudy (legomyego180):

=((x^16))^(1/5)

OpenStudy (anonymous):

exactly but hope you know when there are two exponents like (x^a)^b = x^(a*b) yes ?

OpenStudy (rubylovesjaune):

Yeah

OpenStudy (anonymous):

so than how will be ?

OpenStudy (rubylovesjaune):

x(^5 (square root) x^3?

OpenStudy (anonymous):

\[\sqrt[5]{x^(16)}\]

OpenStudy (anonymous):

= (x^16)^(1/5) = (x^(15/5) *x^(1/5) ) = x^3 *x^(1/5) you have thought it the same ?

OpenStudy (rubylovesjaune):

Yes

OpenStudy (anonymous):

\[x^3*\sqrt[5]{x}\]

OpenStudy (anonymous):

in this way - right ?

OpenStudy (rubylovesjaune):

Yes

OpenStudy (rubylovesjaune):

Thanks you for your help!

OpenStudy (anonymous):

ok was my pleasure good luck bye bye

OpenStudy (rubylovesjaune):

Bye